A puck is moving on an air hockey table. Relative to an xy coordinate system at time t = 0 s, the xcomponents of the puck's initial velocity and acceleration are v0x = +1.2 m/s and ax = +6.6 m/s2. The ycomponents of the puck's initial velocity and acceleration are v0y = +6.5 m/s and ay = -3.3 m/s2. Find (a) the magnitude v and (b) the direction θ of the puck's velocity at a time of t = 0.50 s. Specify the direction relative to the +x axis


1
Expert's answer
2021-01-28T20:11:03-0500

The x componet of the velocity after "t = 0.5s" is:


"v_x = v_{0x} + a_xt\\
v_x =1.2m/s + 6.6m/s^2\cdot 0.5 = 4.5m/s"

The y componet of the velocity after "t = 0.5s" is:


"v_y = v_{0y} + a_yt\\
v_y =6.5m/s -3.3m/s^2\cdot 0.5 =4.85m/s"

The magnitude of the velocity is then (by Phyphagorean's theorem):


"v = \sqrt{v_{x}^2+v_{y}^2}\\
v = \sqrt{4.5^2+4.85^2} \approx 6.62m/s"

The angle relative to the positive x axis is:


"\theta = \arctan\left(\dfrac{v_y}{v_x} \right)\\
\theta = \arctan\left(\dfrac{4.85}{4.5} \right) \approx 47.1\degree"

Answer. a) 6.62m/s, b) 47.1 degree.